15:10 MP entered and check the water pressure OK.
15:10 DSSSDs biased - OK.
The plan from https://elog.ph.ed.ac.uk/DESPEC/719
- Energy distribution in x and y for ADC items and cluster
- The cluster size from conversion electrons
- ADC offsets with the 207Bi source after performing a pulser walkthrough
- Threshold test /w a 207Bi source - varying the threshold as while the source is in place and seeing when the 207Bi peaks are cut off.
- Background run without the 207Bi source (after the above tests)
- Possible test with a 152Eu source as well to measure conversion electrons from Eu could also be an option but we would need to discuss how to do this since all the 152Eu source are closed. We could possibly measure the compton edge from this
data.
Currently the n+n ohmic channels are too noisy to be used for the test. So we will disable their readout for the data taking.
All system wide checks passed. ADC calibration not (but we don't care since we aren't looking at waveforms... for now...).
TEMPS OK. Attachment 2.
We loaded the settings from 2025Oct28-14.55.15 which has blanket thresholds of 0x64 (1 MeV) for all the ASICs.
The pulser peaks seem to have gotten wider since yesterday (75 channels). We have now measured a peak width of 167 channels in aida14 (!) (attachment 3). This was only after tightening the grounding of the setup and disabling the y FEEs. Rates in the ohmic FEEs appear to have decreased by a whole order of magnitude excl. aida02 (see attachment 4).
Currently, we have the ohmic FEE64s disabled aida02,04,06,08, in the discriminator channel mask value set 0x0 ---> 0xfffffff.
15:47 We lower the thresholds to 200 keV (2025Oct29-09.55.31)to see how the system responds and what the rates are.
15:54 We see a 207Bi spectrum cleanly ontop of the FEE64s over which the source is place (aida01,aida14,aida03,aida07). Attachment 6. (For comparison the 207Bi spectrum obtained by Oscar Hall in the AIDA NIM paper Attachment 7).
- A quick observation is that the low energy peak around 400 keV is slightly cut off in some of the FEEs (this could suggest that the operating threshold might be slightly different in some of the FEEs).
16:00 Setting up the merger. EXPERIMENTS/AIDA/2025Oct29-16.11.06 setting made for the discriminators mask setting for the y FEEs disabled.
To the reader of this elog: By now it has taken over an hour to start up the DAQ... We give up and will try tomorrow morning |